Our website employs cookies...

Cookies that we use help us to improve your experience. We do not employ cookies that track you beyond our website. Please read our Privacy Policy for more detail.

UK BASED cloud Backup

Cloud Backup . Fully Managed . UK Telephone Support

01689 661030


Free Trial
Status Date
A small number of users may have experienced connectivity issues overnight between 7pm and 3am. This was due to transit issues between data centers. If you have received a missed backup report then please re-run your backup this morning. 22/03/2018 08:38 AM

Due to the severe weather we have implemented our 'Severe weather policy' and for the safety of all staff, we will all be working from home today and possibly for the rest of the week.

All our normal forms of communication will still work, but if you have problems getting through via the phone, please leave a message and that message will be emailed to us and we will respond.

28/02/2018 06:46 AM
20/02/2018 18:55 - Update to connectivity issues seen at City Reach DC - Since changing from Cisco Network to the Juniper Network at City Reach Data Centre, we have now experienced 4 hours of connectivity without any packet loss. We don't expect any further issues. 20/02/2018 18:57 PM

At 11:30 today, the troublesome connectivity will be moved from the old Cisco Network to Juniper Network to join our other, more reliable transit connection. This will take half of our network completely offline for 5 - 10 minutes.

20/02/2018 11:31 AM

As of 8pm on 19th, we have again been experiencing intermittent connectivity issues that may have affected a number of users. The problem is again with the DC and their last report as of 10pm last night wrote:-

“Please be informed that due to tech error on one of our core switch we have activated your backup link as a primary to avoid any outage. However, due to congestion on this link you are facing intermittent connectivity. Our core engineers are currently working towards a fix.

This is impacting more all IP transit customer, we apologize for any inconvenience this may causing you.”

We have chased again this morning as the issue is still there, but very intermittent. Updates will follow.

20/02/2018 09:32 AM
16/02/2018 10:40am - All systems back online. Data centre investigation: 'Root cause has been identified as an issue with one of our core devices supervisor cards.' 16/02/2018 11:12 AM
16/02/2018 3am - Parts of our network are still offline due to a data centre fault this morning. Further updates soon. 16/02/2018 08:44 AM
13/02/2018 at 20:34pm - Some users may have encountered problems with backup services. A power bar failed and although all cabinets have dual power feeds it seems that the APC switch failed also, probably due to a power surge. All systems were reported fully operational by 10pm. Our data centre is investigating the incident. We apologise for any inconvenience. 14/02/2018 09:33 AM
Planned Maintenance 06/02/2018 expected between 11am to 1pm.
Essential firmware upgrade on firewalls.
Outage expected = less than 30 minutes in 3 steps. 10 minuets each
Will affect all services including phones
06/02/2018 11:28 AM
We would like to inform you of some essential maintenance work that is going to take place on Sunday 24th September 2017 at our London Data Centre between the hours of 3-6am. We will update the status once all work has been completed 24/09/2017 14:38 PM
Website and Partner Portal are now Online 24/09/2017 14:37 PM
Update - All services to the office have been restored and have been stable for 2 hours. 18/07/2017 13:22 PM
Update - All London Servers remain online and working without any issues. However, our office is currently offline so we have switched to our DR backup plan for the office. You should still be able to contact us in the usual ways, but if you're having problems getting through on the phones, please try our online chat found at the bottom of every page of our website, or email support@safedatastorage.co.uk. We will update our status shortly. 18/07/2017 09:43 AM
Around 3:30 this afternoon, some users may have experienced connectivity issues to various hosts within our network. We were made aware that the data centre reported connectivity issues due to a suspected power issue. Although the power issue was resolved within minutes, our connectivity suffered until around 6pm. A temporary fix has now been put in place, but the root problem has yet to be identified. We will update this content once we have an update. The only service this temporary fix is affecting is to our Bournemouth replication site. 18/07/2017 09:41 AM
No further latency or time out issues have been observed. 21/04/2017 05:09 AM
Latency and time out issues have been observed on two of our backup servers over the past evening. Engineers have investigated and resolved a connectivity issues with one of the London SANS. We will continue to monitor for the next 24 hours. 20/04/2017 14:27 PM
No known issues issues on our network 07/12/2016 17:06 PM
Reseller Portal now back online 07/12/2016 17:05 PM
Reseller portal currently offline due to Maintenance 07/12/2016 17:04 PM